# Gibbs states and Brownian models for coexisting haze and cloud droplets

### v1

@@ 1. Case I: $\lambda>\lambda_{K}\rightarrow$ mean-dominated activation
  2. Case II: $\lambda_{c}<\lambda<\lambda_{K}\rightarrow$ fluctuation influenced activation
  3. Case III: $\lambda<1\rightarrow$ fluctuation dominated activation
@@
The parameter $\lambda_{K}$ is the critical Köhler threshold, which is derived to be $\lambda\approx 1.00108$, for the type of aerosols considered in Prabhakaran et al. (2020). The steady-state supersaturation value in each case is also derived from Prabhakaran et al. (2020). Their theoretical calculations reveal indeed that Case I corresponds to $\lambda=1\%$. While the value for Case II is not explicitly provided, it can be inferred from our framework that $\lambda$ is close to $\lambda_{h}$ from below since a bimodal DSD is observed, with high peaks in the haze and activated domain: $\lambda=0.1\%$ is taken for this case. Finally, Case III is a subsaturated regime ($\lambda<0$) and as such we choose $\lambda$ to be given by the Köhler’s value $f(X)$ at the haze peak’s location $X=X_{h}$, this gives $\lambda=-1\%$.
@@
With these parameter values for $\lambda$, the parameter $\beta$ can be estimated for Cases I and II according to Eq. (16), in which $X_{c}$ is chosen to be the experimental DSD peak’s location that corresponds to activated droplets. For Case I, the modal activated diameter is $d_{c}=18.109\mu\mathrm{m}$, giving $X_{c}=1.025\mathrm{s}$ and $\beta=9.7\cdot 10^{-3}\mathrm{s}^{-1/2}$. In Case II, the modal activated diameter is $d_{c}=9.141\mu\mathrm{m}$, giving $X_{c}=0.2611\mathrm{s}$ and $\beta=1.4\cdot 10^{-3}\mathrm{s}^{-1/2}$. These modal values are obtained from the data of Prabhakaran et al. (2020). For Case III, though, since the system is subsaturated and haze particles are more numerous—see Table 1 of Prabhakaran et al. (2020)—the term $\beta$ is practically zero. The corresponding original Köhler and M-Köhler curves are shown in Fig. 6a) and b), as dashed and solid curves, respectively, and shown as functions of the diameter $d=2\sqrt{2DX}$. For Case III, because there is not sink term, only the original Köhler curve is shown.
@@
Once the parameters for the relative disturbances, $\sigma_{1}$, $\sigma_{2}$ and $\kappa$ are chosen as listed in Table 2, the corresponding Gibbs states $\rho_{\lambda}$ given by Eq. (9) exhibit striking skills in reproducing the empirical DSDs in each of the three Pi-chamber experiments taken from Prabhakaran et al. (2020). The results are shown in the Figs. 6d), e) and f). The shapes of the Gibbs states are highly correlated to those of the transformed potentials shown in the insets of Fig. 6, obtained by the Lamperti transformation; see Methods. We observe in particular that the potential wells’ locations match the Gibbs states modes’ locations. Recall that the latter transformation allows for identifying the effective potential for the stochastic model (15) subject to size-dependent stochastic disturbances such as given by (LABEL:eta_term_chamber). With this effective potential, Eq. (15) can be interpreted as a Brownian particle evolving within that potential whose fluctuations are just driven by a white noise, independently of the size of the particle.
@@
As sanity check, one observes that the experimental DSDs’ modes are well predicted by Eq. (11), as shown in Fig. 7 of Noise-induced metastability, especially in Cases I and II, where the classical Köhler theory is unable to predict an activated mode, since $\lambda>\lambda_{K}$. Noteworthy is the creation of a metastable haze-state in Case I resulting from the interaction between the noise term $\sigma(X)\dot{W}_{t}$ and the nonlinear drift terms $-f(X)-\beta X^{1/2}$ as revealed by the transformed potential shown in the inset of Figs. 6d); see Noise-induced metastability in Methods.
@@
Table 2: Supersaturation $\lambda$ and parameter values of the sink term and state-dependent noise $\sigma(X)$ in Eq. (LABEL:eta_term_chamber).
@@ DISCUSSION
Köhler theory stands in the foundation of cloud physics. For a given dry aerosol size and composition, it predicts the particle’s state, haze, or activated droplet as a function of the particle size and the supersaturation. The Köhler curve (Fig. 1a), has two fixed-point regimes bounded by the curve’s mode. The curve left to the mode describes a stable fixed point, whereas any particle that resides on the curve right to the mode is on an unstable fixed point such that a perturbation to the right (or upward) will shift the particle to a continuous growth by diffusion and a perturbation to the left (or downward) will send the particle to the haze state. Warm and small clouds that are driven by weak temperature or humidity perturbations Eytan et al. (2020); Altaratz et al. (2021); Hirsch et al. (2017) are superabundant and have an important radiative effect. However, due to their sizes, weak optical signature, and short lifetimes, these clouds are mostly overlooked, and their properties are poorly understood. Such clouds often exist near the transition between haze to cloud Koren et al. (2009) a regime that can be studied in cloud chambers Prabhakaran et al. (2020); Chandrakar et al. (2016).
@@
Here, we address these clouds by extending Köhler theory to population of droplets embedded in turbulent environments. While activation is explained by Köhler’s theory, the reverse process, i.e. deactivation, is dynamically different. It has been documented in previous works Arabas and Shima (2017); Korolev and Mazin (2003) and in the present study that deactivation is intimately tied to the hysteresis phenomenon through which deactivation occurs at a smaller supersaturation compared to the critical Köhler value $\lambda_{K}$. The addition of a sink term in the condensational growth equation— see Eq. (3)— yields hysteresis directly from an analytical perspective: in order to deactivate a family of monodisperse cloud droplets, the supersaturation needs to be brought strictly below the critical value $\lambda_{K}$ predicted by Köhler, see Figs. 1b). The origin of this behavior is the presence of another saddle-node bifurcation that results from the bending of the M-Köhler curve introduced in Multistable Köhler curves This bending translates further into the appearance of another stable equilibrium on the activated droplet spectrum as a balance between condensational instability and sink factors, like supersaturation consumption or particle removal. We refer to the section Multistability in Methods for conditions on the sink function $g$ ensuring for such a bending to occur, and thus the creation of an extra saddle-node bifurcation in Eq. (3).
The introduction of Brownian noise in the condensational growth equation parameterizes the influence of supersaturation fluctuations and gives analytical formulas for computing DSDs, here identified as Gibbs states; see Eq. (9). Such probability distribution is a general formulation of DSDs found in the literature McGraw and Liu (2006); Chandrakar et al. (2020); Krueger (2020), with the addition of size dependent fluctuations. The resulting DSDs are classified into three types according to the mean supersaturation parameter $\lambda$. The first corresponds to the case where supersaturation fluctuations do not have the strength to activate droplets for a prolonged period of time and, hence, the size distribution is unimodal around the haze state. In the second case, activation is allowed and if the threshold $\lambda_{h}$ is surpassed, haze particles cease to be stable and a single DSD mode is observed peaking over the activated sizes. Consequently, even if Köhler’s threshold supersaturation is not attained, there is a finite probability that particles activate and remain so stably. The associated timescales for activation is therefore related to the average time it takes to droplets to gain sufficient energy to surpass the potential barrier of activation— see Fig. 1c)—, and they are given precisely by Kramers’ formula Eq. (12). In fact, when a cloud parcel is adiabatically cooled, supersaturation fluctuates around a slowly increasing mean. Neglecting noisy fluctuations would incur into an overestimation of the superstaruation threshold for activation, as shown in every panel of Fig. 5, where the interplay between noise and activation hysteresis is illustrated. It is realized that on an activation-deactivation cycle, on average, activation occurs at a much lower supersaturations compared to $\lambda_{h}$, and the larger the variance the lower such threshold is.
The confrontation of our population Köhler scheme within our Brownian modeling framework to steady-state cloud experiments conducted in the Pi-chamber has revealed a striking consistency. Particularly at capturing to a high degree of fidelity the complexity of the DSDs’ shapes along with their bimodal features as reported in the experimental study of Prabhakaran et al. (2020), for different saturation regimes. Under such a setting, it remains to investigate the slow adiabatic changes in mean supersaturation. Based on the analytical insights gained on the droplet activation-deactivation hysteresis (see Fig. 5), we foresee that the same hysteresis behavior should also be observed in the experimental framework of Prabhakaran et al. (2020).
Weak clouds in nature are likely to be exposed to aerosols of several sizes and compositions. However, the weak supersaturation may act as a selective filter, making the largest haze bins the leading players, yielding conditions that are not so different from our theoretical setting. Moreover, the additional Brownian noise blurs the particles’ homogeneity due to the monodisperse nature of the initial conditions. Nevertheless, cases with polydisperse particles of different chemical natures and, hence, different Köhler curves are planned to be investigated.
METHODS
@@
Multistability. In order for model (3) to allow for multiple stable equilibria, it is enough if the the M-Köhler curve $f-g$ has at least a local minimum and a local maximum. This property is to guarantee that the zeros of $\lambda-(f-g)$ correspond to at least two stable equilibria and an unstable one, for a range of values of $\lambda$.
@@
We then have the following sufficiency result for the existence of a local maximum and a local minimum:
@@
Proof.
We need to show that the function $\dot{f}-\dot{g}$ has two zeros, and thus it suffices to prove that $\dot{f}-\dot{g}$ has at least two changes of sign. This function reads, for $X>0$, as:
@@
First of all, $\dot{f}$ has a single local and global minimum. Indeed, we see this by examining $\ddot{f}$:
@@
Due to the continuity and smoothness of the functions $f$ and $g$ on $\mathrm{R}^{+}$, the function $\dot{f}-\dot{g}$ must have at least two zeros. ∎
The confining potential. In order for a gradient system to possess a Gibbs state, it is necessary that the associated potential is confining see e.g. (Pavliotis, 2014, §4). We provide here sufficient mathematical conditions by which a sink function $g$ in Eq. (8) ensures the confining property. Generally, it amounts to imposing a sufficiently strong decay rate to the sink term.
@@
Proof.
By the definition of $f$, $g$ and the potential $V_{\lambda}$, we find that:
@@
This proves the confining property. ∎
Hysteresis path algorithm. To produce the hysteresis paths shown in Fig. 5, we solve Eq. (6) for slowly varying values of the supersaturation parameter $\lambda$ and Brownian noise’s parameter $\sigma$ taken to be additive, i.e. $\sigma(X)=\sqrt{2\varepsilon}$. The slow drift of the supersaturation parameter is organized as follows. We start by dividing an interval $[\lambda_{0},\lambda_{N+1}]\supset(\lambda_{c},\lambda_{h})$ uniformly into $N+2$ grid points $\{\lambda_{j}\}_{j=0}^{N+1}$ such that $\lambda_{j}=\lambda_{0}+j\delta\lambda$, with $\delta\lambda=(\lambda_{h}-\lambda_{c})/(N-1)$ and $\lambda_{0}=\lambda_{c}-\delta\lambda/2$. Note that $\lambda_{N+1}>\lambda_{h}$ and $\lambda_{N}<\lambda_{h}$.
To form the lower branch of this hysteresis path, we start with $\lambda=\lambda_{0}$ and from the haze state $X_{h}(\lambda_{0})$, and solve Eq. (6) using an Euler-Maruyama scheme (Kloeden and Platen, , Chapter 9.1) for one timestep; $\mathrm{d}t=1$. Then $\lambda$ is updated to $\lambda_{0}+j\delta\lambda$ and Eq. (6) is iterated over another timestep by still using Euler-Maruyama. The process is repeated $N+1$ times until $\lambda_{N+1}>\lambda_{h}$. The reverse trajectory is obtained by repeating the algorithm, although starting with $\lambda=\lambda_{N+1}$ at $X_{c}(\lambda_{N+1})$. More precisions are given in Algorithm 1 and hereafter.
@@
Figure 7: Noise-induced metastable states. Panels a), b) and c), correspond to the three different supersaturation configurations as indicated in the titles and corresponding to those of Fig.  6. The blue curves show the growth rate function $-V^{\prime}_{\lambda}$ in Eq. (8), such as associated with the M-Köhler curve $f(X)+\beta X^{1/2}$. The red curves show the derivative of the noise covariance after division by a factor four corresponding to the RHS of Eq. (34). The vertical dashed lines indicate the location of the solutions to Eq. (34) which also correspond to the haze, unstable and activated metastable states, denoted by $d_{h},d_{u}$ and $d_{c}$, respectively.
The Lamperti transformation. For scalar stochastic differential model, it is possible to map a stochastic differential equation (SDE) driven by state-dependent noise onto an SDE driven by an additive noise. We recall this basic idea. Consider the following scalar SDE
@@
This is called the Lamperti transformation (Pavliotis, 2014, §3.6).
@@
Noise-induced metastability. By applying (32) to our stochastic droplet model (15) in Pi-chamber, we obtain:
@@
recovering this way Eq. (11). Thus, if one wishes to match the locations of these metastable states with let us say the modes of an empirical distribution (as in PI-Chamber Empirical Distributions vs Gibbs States), one needs to solve Eq. (34). As a consequence, this equation imposes constraints on the choice of the parameters involved in $\sigma$ and $V_{\lambda}$.
@@
The blue and red curves in Fig. 7 correspond to the left-hand-side (LHS) and right-hand-side (RHS) of Eq. (34), respectively, for the three case studies of PI-Chamber Empirical Distributions vs Gibbs States. The intersection of these two curves determines the location of the metastable states indicated by the dashed vertical lines for Cases I, II and III with the parameter configuration shown in Table 2.
Acknowledgements This work has been partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 810370). MSG is grateful to the Feinberg Graduate School for their support through the Dean of Faculty Fellowship.
Author contributions: MDC conceived the presented idea. MSG led the analyses and MDC supported. MSG, MDC, OA, and IK discussed the results and wrote the manuscript. All co-authors provided critical feedback and helped shape the research, analysis and manuscript.
Competing interests: The authors declare no competing interests.
Data availability: All experimental data used in this study are based on the article of Prabhakaran et al. (2020) and are publicly available at: https://digitalcommons.mtu.edu/data-files/3/. The numerical data obtained for this article are available upon request to the authors.

### v2

@@ 1. Case I: $\lambda>\lambda_{K}\rightarrow$ mean-dominated activation
  2. Case II: $\lambda_{c}<\lambda<\lambda_{K}\rightarrow$ fluctuation-influenced activation
  3. Case III: $\lambda<1\rightarrow$ fluctuation-dominated activation
@@
The parameter $\lambda_{K}$ is the critical Köhler threshold, which is derived to be $\lambda\approx 1.00108$, for the type of aerosols considered in [PSK+20]. The steady-state supersaturation value in each case is also derived from [PSK+20]. Their theoretical calculations reveal indeed that Case I corresponds to $\lambda=1\%$. While the value for Case II is not explicitly provided, it can be inferred from our framework that $\lambda$ is close to $\lambda_{h}$ from below since a bimodal DSD is observed, with high peaks in the haze and activated domain: $\lambda=0.1\%$ is taken for this case. Finally, Case III is a subsaturated regime ($\lambda<0$) and as such we choose $\lambda$ to be given by the Köhler’s value $f(X)$ at the haze peak’s location $X=X_{h}$, this gives $\lambda=-1\%$.
@@
With these parameter values for $\lambda$, the parameter $\beta$ can be estimated for Cases I and II according to Eq. (16), in which $X_{c}$ is chosen to be the experimental DSD peak’s location that corresponds to activated droplets. For Case I, the modal activated diameter is $d_{c}=18.109\mu\mathrm{m}$, giving $X_{c}=1.025\mathrm{s}$ and $\beta=9.7\cdot 10^{-3}\mathrm{s}^{-1/2}$. In Case II, the modal activated diameter is $d_{c}=9.141\mu\mathrm{m}$, giving $X_{c}=0.2611\mathrm{s}$ and $\beta=1.4\cdot 10^{-3}\mathrm{s}^{-1/2}$. These modal values are obtained from the data of [PSK+20]. For Case III, though, since the system is subsaturated and haze particles are more numerous—see Table 1 of [PSK+20]—the term $\beta$ is practically zero. The corresponding original Köhler and M-Köhler curves are shown in Fig. 6a) and b), as dashed and solid curves, respectively, and shown as functions of the diameter $d=2\sqrt{2DX}$. For Case III, because there is no sink term, only the original Köhler curve is shown in Fig. 6c).
@@
Once the parameters for the relative disturbances, $\sigma_{1}$, $\sigma_{2}$ and $\kappa$ are chosen as listed in Table 2, the corresponding Gibbs states $\rho_{\lambda}$ given by Eq. (9) exhibit striking skills in reproducing the empirical DSDs in each of the three Pi-chamber experiments taken from [PSK+20]. The results are shown in the Figs. 6d), e) and f). The shapes of the Gibbs states are highly correlated to those of the transformed potentials shown in the insets of Fig. 6, obtained by the Lamperti transformation; see Section "The Lamperti transformation" in Material and Methods. We observe in particular that the potential wells’ locations match the Gibbs states modes’ locations. Recall that the latter transformation allows for identifying the effective potential for the stochastic model (15) subject to size-dependent stochastic disturbances such as given by (LABEL:eta_term_chamber). With this effective potential, Eq. (15) can be interpreted as a Brownian particle evolving within that potential whose fluctuations are just driven by a white noise, independently of the size of the particle.
@@
The theory allows us to make striking predictions. For instance, the experimental DSDs’ modes are well predicted by Eq. (11), as shown in Fig. 7 of Section "Noise-induced metastability" in Material and Methods, especially in Cases I and II, where the classical Köhler theory is unable to predict an activated mode, since $\lambda>\lambda_{K}$. Noteworthy is the creation of a metastable haze-state in Case I resulting from the interaction between the noise term $\sigma(X)\dot{W}_{t}$ and the nonlinear drift terms $-f(X)-\beta X^{1/2}$ as revealed by the transformed potential shown in the inset of Figs. 6d); see Section "Noise-induced metastability" in Material and Methods.
@@
Table 2. Supersaturation values, sink term’s and state-dependent noise’s parameters.
@@ Discussion
Köhler theory is at the basis of cloud physics. For a given dry aerosol size and composition, it predicts the particle’s state, haze, or activated droplet as a function of the particle size and the supersaturation. The Köhler curve (Fig. 1a), has two fixed-point regimes bounded by the curve’s mode. The curve left to the mode describes a stable fixed point, whereas any particle that resides on the curve right to the mode is on an unstable fixed point such that a perturbation to the right (or upward) will shift the particle to a continuous growth by diffusion and a perturbation to the left (or downward) will send the particle to the haze state. Warm and small clouds that are driven by weak temperature or humidity perturbations [EKA+20, AKA+21, HKA+17] are abundant and have an important radiative effect. However, due to their sizes, weak optical signature, and short lifetimes, these clouds are mostly overlooked, and their properties are poorly understood. Such clouds often exist near the transition between haze to cloud [KFJA09] a regime that can be studied in cloud chambers [PSK+20, CCC+16].
@@
Here, we address these clouds by extending Köhler theory to a population of droplets embedded in turbulent environments. While activation is explained by Köhler’s theory, the reverse process, i.e. deactivation, is dynamically different. It has been documented in previous works [AS17, KM03] and in the present study that deactivation is intimately tied to the hysteresis phenomenon through which deactivation occurs at a smaller supersaturation compared to the critical Köhler value $\lambda_{K}$. The addition of a sink term in the condensational growth equation — see Eq. (3) — yields hysteresis directly from an analytical perspective: in order to deactivate a family of monodisperse cloud droplets, the supersaturation needs to be brought strictly below the critical value $\lambda_{K}$ predicted by Köhler. The origin of this behavior is the presence of another saddle-node bifurcation that results from the bending of the M-Köhler curve introduced in Section "Multistable Köhler curves." This bending translates further into the appearance of another stable equilibrium on the activated droplet spectrum as a balance between condensational instability and sink factors, like supersaturation consumption or particle removal.
The introduction of Brownian noise in the condensational growth equation parameterizes the influence of supersaturation fluctuations and gives analytical formulas for computing DSDs, here identified as Gibbs states; see Eq. (9). Such probability distributions generalize earlier formulas for DSDs found in the literature such as Weibull distributions [ML06, CSY+20, Kru20].
The resulting Droplet Size Distributions (DSDs) are categorized into three types based on the mean supersaturation parameter $\lambda$. In the first type, fluctuations in supersaturation are insufficient to sustain droplet activation for extended periods, leading to a unimodal DSD centered around the haze state. In the second type, activation is possible, and when the threshold $\lambda_{h}$ is exceeded, haze particles become unstable. This results in a single DSD mode peaking at activated sizes. Notably, even if Köhler’s threshold supersaturation is not reached, there is a finite probability of particle activation. The associated activation timescales are linked to the average time required for droplets to overcome the activation energy barrier (Fig. 1c), as accurately described by Kramers’ formula (Eq. (12)).
When a cloud parcel undergoes adiabatic cooling, supersaturation fluctuates around a gradually increasing mean. Ignoring turbulent fluctuations can lead to an overestimation of the activation threshold, as illustrated in Figure 5. This figure depicts various hysteresis loops where activation occurs at even lower supersaturation values than $\lambda_{h}$, especially when the noise variance is large.
Our stochastic M-Köhler model has demonstrated remarkable agreement with steady-state cloud experiments conducted in the Pi-chamber. Notably, it accurately captures the intricate DSD shapes, including bimodal features, as observed in the experimental study by [PSK+20] across various saturation regimes. To further explore this consistency, future investigations should focus on slow adiabatic changes in mean supersaturation. Based on the analytical understanding of droplet activation-deactivation hysteresis (Fig. 5), we anticipate that similar hysteresis behavior will be observed in the experimental setup of [PSK+20].
In natural environments, weak clouds often encounter aerosols of varying sizes and compositions. Nevertheless, the low supersaturation levels may act as a selective filter, favoring the activation of larger haze particles. This can create conditions that are not substantially different from our theoretical framework. Additionally, the presence of Brownian motion can introduce variability into the particle distribution, even with initially monodisperse conditions. Future studies will explore scenarios involving polydisperse particles with diverse chemical properties and varying Köhler curves.
Acknowledgments
The insights discussed in this work have substantially benefited from the reviewers’ constructive comments, and we express our deep gratitude to them.
Funding:
This work has been partially supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement no. 810370). MSG is grateful to the Feinberg Graduate School for their support through the Dean of Faculty Fellowship.
Author contributions:
MDC conceived the presented idea. MSG led the analyses and MDC supported. MSG, MDC, and IK discussed the results and wrote the manuscript. All co-authors provided critical feedback and helped shape the research, analysis and manuscript.
Competing interests:
The authors declare no competing interests.
Data and materials availability:
All experimental data used in this study are based on the article of [PSK+20] and are publicly available at: https://digitalcommons.mtu.edu/data-files/3/. All data needed to evaluate the conclusions in the paper are present in the paper and/or the Supplementary Materials.
Material and Methods
Multistability
@@ Acknowledgments
To ensure that the model defined by Equation (3) exhibits multiple stable equilibria, it suffices for the M-Köhler curve $f-g$ to possess at least one local minimum and one local maximum. This condition guarantees that the roots of the equation $\lambda-(f-g)=0$ correspond to at least two stable equilibria and one unstable equilibrium within a certain range of $\lambda$ values.
@@
We then have the following sufficiency conditions for the existence of a local maximum and a local minimum:
@@
Proof. To demonstrate the existence of two zeros for the function $\dot{f}-\dot{g}$, it is sufficient to prove that it undergoes at least two sign changes. For $X>0$, this function is given by:
@@
First of all, $\dot{f}$ has a single local and global minimum. Indeed, this can be observed by examining $\ddot{f}$:
@@
Due to the continuity and smoothness of the functions $f$ and $g$ on $\mathrm{R}^{+}$, the function $\dot{f}-\dot{g}$ must have at least two zeros.
The confining potential
In order for a gradient system to possess a Gibbs state, it is necessary that the associated potential is confining see e.g. Chapter 4 in [Pav14]. We provide here sufficient mathematical conditions by which a sink function $g$ in Eq. (8) ensures the confining property. Generally, it amounts to imposing a sufficiently strong decay rate to the sink term.
@@
Proof. By the definition of $f$, $g$ and the potential $V_{\lambda}$, we find that:
@@
This proves the confining property.
Hysteresis path algorithm
To produce the hysteresis paths shown in Fig. 5, we solve Eq. (6) for slowly varying values of the supersaturation parameter $\lambda$ and Brownian noise’s parameter $\sigma$ taken to be additive, i.e. $\sigma(X)=\sqrt{2\varepsilon}$. The slow drift of the supersaturation parameter is organized as follows. We start by dividing an interval $[\lambda_{0},\lambda_{N+1}]\supset(\lambda_{c},\lambda_{h})$ uniformly into $N+2$ grid points $\{\lambda_{j}\}_{j=0}^{N+1}$ such that $\lambda_{j}=\lambda_{0}+j\delta\lambda$, with $\delta\lambda=(\lambda_{h}-\lambda_{c})/(N-1)$ and $\lambda_{0}=\lambda_{c}-\delta\lambda/2$. Note that $\lambda_{N+1}>\lambda_{h}$ and $\lambda_{N}<\lambda_{h}$.
To form the lower branch of this hysteresis path, we start with $\lambda=\lambda_{0}$ and from the haze state $X_{h}(\lambda_{0})$, and solve Eq. (6) using an Euler-Maruyama scheme with time-step $\delta t=10^{-2}\mathrm{s}$; see Chapter 9.1 in [KP92].
Then $\lambda$ is updated to $\lambda_{0}+j\delta\lambda$ and Eq. (6) is iterated over another timestep by still using Euler-Maruyama. The process is repeated $N+1$ times until $\lambda_{N+1}>\lambda_{h}$. The reverse trajectory is obtained by repeating the algorithm, although starting with $\lambda=\lambda_{N+1}$ at $X_{c}(\lambda_{N+1})$. More precisions are given in Algorithm 1 and hereafter.
@@
Figure 7. Noise-induced metastable states. Panels a), b) and c), correspond to the three different supersaturation configurations as indicated in the titles and corresponding to those of Fig.  6. The blue curves show the growth rate function $-V^{\prime}_{\lambda}$ in Eq. (8), such as associated with the M-Köhler curve $f(X)+\beta X^{1/2}$. The red curves show the derivative of the noise covariance after division by a factor four corresponding to the RHS of Eq. (34). The vertical dashed lines indicate the location of the solutions to Eq. (34) which also correspond to the haze, unstable and activated metastable states, denoted by $d_{h},d_{u}$ and $d_{c}$, respectively.
The Lamperti transformation
For scalar stochastic differential model, it is possible to map a stochastic differential equation (SDE) driven by state-dependent noise onto an SDE driven by an additive noise. We recall this basic idea. Consider the following scalar SDE
@@
This is called the Lamperti transformation; see Chapter 3.6 in [Pav14].
@@
Noise-induced metastability
By applying (32) to our stochastic droplet model (15) in Pi-chamber, we obtain:
@@
recovering this way Eq. (11). Thus, if one wishes to match the locations of these metastable states with let us say the modes of an empirical distribution (as in Section "Pi-Chamber empirical distributions vs Gibbs states"), one needs to solve Eq. (34). As a consequence, this equation imposes constraints on the choice of the parameters involved in $\sigma$ and $V_{\lambda}$.
@@
The blue and red curves in Fig. 7 correspond to the left-hand-side (LHS) and right-hand-side (RHS) of Eq. (34), respectively, for the three case studies of Section "Pi-Chamber empirical distributions vs Gibbs states." The intersection of these two curves determines the location of the metastable states indicated by the dashed vertical lines for Cases I, II and III with the parameter configuration shown in Table 2.